# references.bib

### v1

@@ 1 Introduction
Recent foundation models, such as OpenAI’s o1 [openaiteamIntroducingOpenAIO12024] and DeepSeek R1 [guoDeepSeekR1IncentivizingReasoning2025], have shown that incorporating explicit reasoning phases before generating answers significantly improves performance across tasks. Despite these advances, audio-language models have rarely implemented such structured reasoning approaches. Although general foundation models increasingly separate generation into explicit “thinking” and “answering” stages, audio understanding models still rely largely on end-to-end architectures that directly map audio into text without intermediate reasoning steps. This gap represents a promising opportunity to improve audio understanding capabilities through more deliberate reasoning mechanisms.
Several techniques have emerged to extend test-time reasoning in a controlled and scalable manner. One such method is budget forcing, which inserts “wait” tokens during generation to lengthen the model’s reasoning process [muennighoffS1SimpleTesttime2025]. Other approaches rely on reinforcement learning to manage a model’s thinking budget—the computational effort allocated during inference. For instance, Length-Controlled Policy Optimization (LCPO) [aggarwalL1ControllingHow2025] jointly optimizes for accuracy and reasoning length, while cosine-shaped reward functions have been used to penalize overly long yet incorrect outputs [dangReinforcementLearningReasoning2025]. Notably, scaling test-time compute has shown significant benefits, with smaller models like LLaMA 3B outperforming much larger ones, such as LLaMA 70B, when allowed more extensive reasoning during inference [beechingScalingTesttimeCompute].
@@
A parallel challenge in audio-language modeling is the limited diversity of training data. The field has become highly compartmentalized, with most models trained on overlapping datasets primarily sourced from AudioSet [gemmekeAudioSetOntology2017] and Freesound [fontcorberaFreesoundTechnicalDemo2013]. This redundancy is reflected in widely used benchmarks such as AudioCaps [kimAudioCapsGeneratingCaptions2019], WavCaps [meiWavCapsChatGPTassistedWeaklylabelled2024], and various derivatives. For example, studies report that 17.6% of AudioCaps overlap with WavCaps, while Clotho [drossosClothoAudioCaptioning2020] exhibits even higher overlaps - up to 89% - with other datasets [labbeCoNeTTEEfficientAudio2024]. Such cross-contamination contributes to a homogenized dataset landscape, limiting the generalization and robustness of audio-language models [wijngaardAudioLanguageDatasetsScenes2025].
To address the dual challenges of limited reasoning in audio-language models and the lack of diverse training data, we introduce AudSemThinker (Audio Semantics Thinker), a model that grounds its reasoning process in a structured framework of auditory semantics inspired by human cognition. Specifically, the model analyzes audio in terms of functional components, including sound-generating agents (who), physical sound sources (what), generation mechanisms (how), and contextual cues (when/where). This framework draws on theoretical foundations [gaverWhatWorldWe1993] and experimental research [gygiSimilarityCategorizationEnvironmental2007, bonesSoundCategoriesCategory2018] that investigate the perceptual and cognitive mechanisms driving human perception of complex sounds [giordanoWhatWeMean2022]. To support the development of AudSemThinker, we also construct AudSem, a new dataset synthesized from YouTube closed captions, designed to enhance training diversity and reduce overlap with existing resources. Using the AudSem dataset, we train AudSemThinker using two paradigms, Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO), and evaluate its performance on a range of established benchmarks, spanning both general audio foundation tasks and audio-based dialogue.
@@ 2.1 Audio Datasets with Synthetically Generated Captions
Early audio-language datasets primarily relied on manual annotation or existing metadata. For example, AudioCaps [kimAudioCapsGeneratingCaptions2019] employed Amazon Mechanical Turk workers to generate captions for AudioSet clips. Subsequently, the SAM-S dataset [hebbarDatasetAudioVisualSound2023] introduced an innovative approach by extracting sound-specific data from closed caption transcripts of Hollywood movies, yielding 116,000 captions. A significant shift occurred with the advent of AI-assisted dataset creation for audio models. In the audio domain, 80% of new datasets created in the first half of 2024 incorporated LLM-generated captions, a marked increase from 69.2% in 2023 [wijngaardAudioLanguageDatasetsScenes2025]. One notable example is Auto-ACD [sunAutoACDLargescaleDataset2024], which employed multiple AI models to process both audio and visual information from VGGSound and AudioSet. The dataset creation pipeline integrated seven different models, including BLIP-2 [liBLIP2BootstrappingLanguageImage2023] for image captioning and the CLIP model [radfordLearningTransferableVisual2021] for categorization. Other examples of AI-assisted dataset creation for audio models include Sound-VECaps [yuanSoundVECapsImprovingAudio2024], which employed CogVLM [wangCogVLMVisualExpert2024] for visual captioning and LLaMA 3 [llama3teamLlama3Herd2024] for generating audio descriptions. This approach produced two complementary dataset variants: Sound-VECapsF, which incorporates detailed visual features, and Sound-VECapsA, which focuses solely on audible content. Building on this, AudioSetCaps [baiAudioSetCapsEnrichedAudio2024] introduced a more refined three-stage pipeline that combines large audio-language models for content extraction, LLMs for caption generation, and CLAP-based refinement [wuLargeScaleContrastiveLanguageAudio2023], resulting in 1.9 million high-quality audio-caption pairs.
@@
WavCaps [meiWavCapsChatGPTassistedWeaklylabelled2024] Metadata 1 44.87
AudioSetCaps [baiAudioSetCapsEnrichedAudio2024] Labels 4 179.20
AF-AudioSet [kongImprovingTextToAudioModels2024] - 1 143.57
Auto-ACD [sunAutoACDLargescaleDataset2024] Labels 7 103.87
Sound-VECapsF [yuanSoundVECapsImprovingAudio2024] Labels 3 180.72
AudSem (ours) CC 9 852.63
@@ 2.2 Audio Models with Reasoning
Several recent efforts have focused on integrating reasoning capabilities into audio-language models. One example is Mellow [deshmukhMellowSmallAudio2025], a lightweight audio-language model that demonstrated strong reasoning abilities. Despite having only 167 million parameters and being trained on 1.24 million examples, Mellow outperformed larger state-of-the-art models across various domains. Another prominent model, Audio-Reasoner [xieAudioReasonerImprovingReasoning2025], was specifically designed for deep reasoning in audio tasks. It introduced a structured reasoning process using a large-scale dataset (CoTA) and a multi-phase “thinking” architecture consisting of planning, captioning, reasoning, and summarization before generating its final response. Audio-CoT [maAudioCoTExploringChainofThought2025] explored the integration of Chain-of-Thought (CoT) prompting within large audio-language models, revealing a positive correlation between the length of reasoning paths and model accuracy.
@@ 3 AudSem Dataset
The initial dataset creation phase started with a vast collection of YouTube subtitles, filtering for bracketed SDH captions likely representing sound descriptions. The YouTube subtitles were further refined using both BERT [devlinBERTPretrainingDeep2019] classifier and a Mixtral [jiangMixtralExperts2024] decoder to identify genuine audio events, resulting in 9.5 million potential captions. The corresponding video segments were downloaded, and audio/video streams were standardized (see Appendix C.1 for details).
The following stage involved analyzing the extracted audio-visual segments using multiple specialized models. Audio analysis employed an ensemble including Qwen2Audio-7B [chuQwen2AudioTechnicalReport2024a], BEATs [chenBEATsAudioPretraining2023], AST [gongASTAudioSpectrogram2021], CoNeTTE [labbeCoNeTTEEfficientAudio2024], and LP-MusicCaps Model [dohLPMusicCapsLLMBasedPseudo2023]. Visual analysis used BLIP [liBLIPBootstrappingLanguageImage2022], CLIP [radfordLearningTransferableVisual2021], RT-DETR [zhaoDETRsBeatYOLOs2024a], and Places365 [zhouLearningDeepFeatures2014] on sampled frames, filtering black frames and deduplicating results. Video analysis utilized LLaVA-Video-7B [zhangVideoInstructionTuning2024] to capture temporal context from uniformly sampled frames. This stage yielded an initial dataset of 5,332,211 samples, which then underwent multiple filtering steps to ensure quality. Outliers were removed if the cosine distance between their CLAP embedding and the average embedding exceeded 0.9 (for both audio and text). Audio samples shorter than three seconds were also excluded. A final filtering step ensured that the cosine similarity between generated audio captions and original YouTube closed captions was at least 0.5, resulting in 213,908 high-quality samples (see Appendix C.2 for details).
The final stage utilized the Qwen2.5-72B-Instruct model [yangQwen25TechnicalReport2025] to synthesize the final sound description from the processed multi-modal features. This involved a structured generation approach with enforced JSON output, implementing a thinking phase, an optional semantic descriptor phase, and an answer phase. A separate judging model validated outputs for quality and adherence to guidelines, with regeneration attempts for failed outputs. From the 213k processed examples, four types of tasks were generated (captioning, multiple-choice QA, open-ended QA, creative writing), resulting in approximately 800k-900k final examples depending on the inclusion of semantic descriptors (see Appendix C.3 for details). When comparing against other datasets, there was an overlap of only 12 examples with AudioSet, with one example also being in AudioCaps, and 0 in VGGSound.
@@
Both configurations use XML-style tags to distinguish between phases. In the two-phase approach, the model’s output is structured with <thinking> and <answer> tags. The three-phase approach introduces an additional <semantic_elements> section between thinking and answering, allowing the model to explicitly represent key semantic descriptors before generating the final caption. See also Appendix E with Figure 5 for an example of an output from the two-phase approach, and Figure 6 for an example of an output from the three-phase approach.
@@ 3.1 Semantic Descriptors
The core idea behind incorporating semantic descriptors into model reasoning is to capture the detailed aspects of audio that a human listener would naturally perceive (see also [giordanoWhatWeMean2022, wijngaardACESEvaluatingAutomated2023]). Before generating a final caption, the model first engages in a structured “thinking” phase, during which it attempts to identify and describe specific components of the audio signal. This process enables the model to move beyond surface-level recognition toward a more nuanced understanding of what it is “hearing”. The descriptors cover both objective and subjective dimensions of auditory perception. Objective descriptors include elements that produce the sound, such as sound-generating agents and physical sound sources, while subjective descriptors address how the sound is perceived, including auditory attributes and non-auditory sensations. This structured reasoning approach allows the model to decompose complex audio scenes into their core semantic descriptors, resulting in more informative and interpretable outputs. See Table 2 for definitions of the semantic descriptors used in our approach.
@@ 4 AudSemThinker Model
The proposed approach employs the Qwen2.5-Omni [yangQwen25TechnicalReport2025] model with 7B parameters. Qwen2.5-Omni is an end-to-end multimodal model that processes diverse inputs including text, images, audio, and video, and generates both text and speech outputs through its Thinker-Talker architecture. For AudSemThinker, the “Thinker” component, the large language model responsible for text generation, is specifically fine-tuned. This architecture processes multi-modal inputs and generates textual outputs. Fine-tuning occurs exclusively on the AudSem dataset (see Section 3); no other fine-tuning is performed. The core idea involves guiding the model to first engage in a “thinking” process, describe the semantic descriptors of the audio, and then produce a final “answer”. Two primary training paradigms are explored: SFT and GRPO, both using the AudSem dataset, demonstrating the dataset’s effectiveness in both scenarios.
@@ 4.1 Supervised Fine-Tuning (SFT)
Parameter-efficient fine-tuning is used with LoRA [huLoRALowrankAdaptation2022], targeting projection layers. Both models are trained using the AdamW optimizer [loshchilovDecoupledWeightDecay2019a] with a learning rate of 2e-04 for one epoch, employing a scheduler with linear decay and bf16 precision. Training uses a batch size of four on a single H100 GPU, taking approximately 12 hours for the full dataset (AudSemThinker) and six hours for the QA subset (AudSemThinker-QA). A detailed instructional prompt is prepended to the user’s question (see Appendix F.3), where the model is guided on the expected output format and reasoning depth (see Appendix 5.3 for a detailed ablation study).
@@ 4.2 Reinforcement Learning with Group Relative Policy Optimization (GRPO)
Beyond supervised fine-tuning, we investigate the applicability of reinforcement learning to our task and dataset using GRPO [shaoDeepSeekMathPushingLimits2024]. GRPO has shown to be an effective technique in finetuning models and has been used to train state-of-the-art models such as DeepSeek R1 [shaoDeepSeekMathPushingLimits2024] and Qwen2.5 [yangQwen25TechnicalReport2025]. GRPO allows us to explore whether the AudSem dataset is suitable for RL-based finetuning and to specifically test the impact of controlling the model’s “thinking budget” during generation. GRPO with Verifiable Rewards (RLVR) [cobbeTrainingVerifiersSolve2021, suExpandingRLVerifiable2025] is employed, building upon a Qwen2.5-Omni base model.
@@
  • Length Constraint Reward: Specifically targeting the <think> phase, this reward function implements a “thinking budget”. It penalizes deviations from a target length for the reasoning content, using parameters to control the penalty strength and tolerance. We adapt the approach from [aggarwalL1ControllingHow2025] (with $\alpha=0.1,\delta=0.5$) by adding a reward component for being close to the target length (see Appendix D for details). This encourages the model to generate thinking content near a specified target length while penalizing excessive length.
These reward functions produce scalars for each generation, which are weighted and summed. For the group of generations corresponding to each prompt, the mean and standard deviation of these summed rewards are calculated. An advantage (summed reward -- group mean) is computed for each generation, and these advantages are optionally normalized using the group’s standard deviation to produce the final scalar reward signal used for policy updates.
@@
The resulting model, AudSemThinker-QA GRPO, is trained with a target thinking budget of 25 words. Similar to SFT, LoRA [huLoRALowrankAdaptation2022] is used for parameter-efficient tuning. The input prompts for GRPO also include instructions referencing the target thinking length. For training, we use the default GRPO loss type with $\beta=0.01$, generate six completions per prompt ($k=6$), and employ a batch size of two per device with bf16 precision. The model is trained on four H100 GPUs for approximately 10 hours, utilizing DeepSpeed ZeRO-3 [rajbhandariZeROMemoryOptimizations2020] and vLLM [kwonEfficientMemoryManagement2023] for efficient training and inference.

### v2

# AudSemThinker: Enhancing Audio-Language Models through Reasoning over Semantics of Sound

@@ 1 Introduction
Recent foundation models, such as OpenAI’s o1 [49] and DeepSeek R1 [25], have shown that incorporating explicit reasoning phases before generating answers significantly improves performance across tasks. Despite these advances, audio-language models have rarely implemented such structured reasoning approaches. Although general foundation models increasingly separate generation into explicit “thinking” and “answering” stages, audio understanding models still rely largely on end-to-end architectures that directly map audio into text without intermediate reasoning steps. This gap represents a promising opportunity to improve audio understanding capabilities through more deliberate reasoning mechanisms.
Several techniques have emerged to extend test-time reasoning in a controlled and scalable manner. One such method is budget forcing, which inserts “wait” tokens during generation to lengthen the model’s reasoning process [47]. Other approaches rely on reinforcement learning to manage a model’s thinking budget—the computational effort allocated during inference. For instance, Length-Controlled Policy Optimization (LCPO) [2] jointly optimizes for accuracy and reasoning length, while cosine-shaped reward functions have been used to penalize overly long yet incorrect outputs [11]. Notably, scaling test-time compute has shown significant benefits, with smaller models like LLaMA 3B outperforming much larger ones, such as LLaMA 70B, when allowed more extensive reasoning during inference [4].
@@
A parallel challenge in audio-language modeling is the limited diversity of training data. The field has become highly compartmentalized, with most models trained on overlapping datasets primarily sourced from AudioSet [20] and Freesound [17]. This redundancy is reflected in widely used benchmarks such as AudioCaps [33], WavCaps [45], and various derivatives. For example, studies report that 17.6% of AudioCaps overlap with WavCaps, while Clotho [16] exhibits even higher overlaps - up to 89% - with other datasets [37]. Such cross-contamination contributes to a homogenized dataset landscape, limiting the generalization and robustness of audio-language models [63].
To address the dual challenges of limited reasoning in audio-language models and the lack of diverse training data, we introduce AudSemThinker (Audio Semantics Thinker), a model that grounds its reasoning process in a structured framework of auditory semantics inspired by human cognition. Specifically, the model analyzes audio in terms of functional components, including sound-generating agents (who), physical sound sources (what), generation mechanisms (how), and contextual cues (when/where). This framework draws on theoretical foundations [18] and experimental research [26, 5] that investigate the perceptual and cognitive mechanisms driving human perception of complex sounds [23]. To support the development of AudSemThinker, we also construct AudSem, a new dataset synthesized from YouTube closed captions, designed to enhance training diversity and reduce overlap with existing resources. Using the AudSem dataset, we train AudSemThinker using two paradigms, Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO), and evaluate its performance on a range of established benchmarks, spanning both general audio foundation tasks and audio-based dialogue.
@@ 2.1 Audio Datasets with Synthetically Generated Captions
Early audio-language datasets primarily relied on manual annotation or existing metadata. For example, AudioCaps [33] employed Amazon Mechanical Turk workers to generate captions for AudioSet clips. Subsequently, the SAM-S dataset [28] introduced an innovative approach by extracting sound-specific data from closed caption transcripts of Hollywood movies, yielding 116,000 captions. A significant shift occurred with the advent of AI-assisted dataset creation for audio models. In the audio domain, 80% of new datasets created in the first half of 2024 incorporated LLM-generated captions, a marked increase from 69.2% in 2023 [63]. One notable example is Auto-ACD [57], which employed multiple AI models to process both audio and visual information from VGGSound and AudioSet. The dataset creation pipeline integrated seven different models, including BLIP-2 [39] for image captioning and the CLIP model [52] for categorization. Other examples of AI-assisted dataset creation for audio models include Sound-VECaps [68], which employed CogVLM [60] for visual captioning and LLaMA 3 [42] for generating audio descriptions. This approach produced two complementary dataset variants: Sound-VECapsF, which incorporates detailed visual features, and Sound-VECapsA, which focuses solely on audible content. Building on this, AudioSetCaps [3] introduced a more refined three-stage pipeline that combines large audio-language models for content extraction, LLMs for caption generation, and CLAP-based refinement [64], resulting in 1.9 million high-quality audio-caption pairs.
@@
WavCaps [45] Metadata 1 44.87
AudioSetCaps [3] Labels 4 179.20
AF-AudioSet [35] - 1 143.57
Auto-ACD [57] Labels 7 103.87
Sound-VECapsF [68] Labels 3 180.72
AudSem (ours) $C$ 9 852.63
@@ 2.2 Audio Models with Reasoning
Several recent efforts have focused on integrating reasoning capabilities into audio-language models. One example is Mellow [12], a lightweight audio-language model that demonstrated strong reasoning abilities. Despite having only 167 million parameters and being trained on 1.24 million examples, Mellow outperformed larger state-of-the-art models across various domains. Another prominent model, Audio-Reasoner [65], was specifically designed for deep reasoning in audio tasks. It introduced a structured reasoning process using a large-scale dataset (CoTA) and a multi-phase “thinking” architecture consisting of planning, captioning, reasoning, and summarization before generating its final response. Audio-CoT [44] explored the integration of Chain-of-Thought (CoT) prompting within large audio-language models, revealing a positive correlation between the length of reasoning paths and model accuracy.
@@ 3 AudSem Dataset
The initial dataset creation phase started with a vast collection of YouTube subtitles, filtering for bracketed SDH captions likely representing sound descriptions. The YouTube subtitles were further refined using both BERT [13] classifier and a Mixtral [32] decoder to identify genuine audio events, resulting in 9.5 million potential captions. The corresponding video segments were downloaded, and audio/video streams were standardized (see Appendix B.1 for details).
The following stage involved analyzing the extracted audio-visual segments using multiple specialized models. Audio analysis employed an ensemble including Qwen2Audio-7B [8], BEATs [6], AST [24], CoNeTTE [37], and LP-MusicCaps Model [14]. Visual analysis used BLIP [40], CLIP [52], RT-DETR [70], and Places365 [71] on sampled frames, filtering black frames and deduplicating results. Video analysis utilized LLaVA-Video-7B [69] to capture temporal context from uniformly sampled frames. This stage yielded an initial dataset of 5,332,211 samples, which then underwent multiple filtering steps to ensure quality. Outliers were removed if the cosine distance between their CLAP embedding and the average embedding exceeded 0.9 (for both audio and text). Audio samples shorter than three seconds were also excluded. A final filtering step ensured that the cosine similarity between generated audio captions and original YouTube closed captions was at least 0.5, resulting in 213,908 high-quality samples (see Appendix B.2 for details).
The final stage utilized the Qwen2.5-72B-Instruct model [67] to synthesize the final sound description from the processed multi-modal features. This involved a structured generation approach with enforced JSON output, implementing a thinking phase, an optional semantic descriptor phase, and an answer phase. A separate judging model validated outputs for quality and adherence to guidelines, with regeneration attempts for failed outputs. From the 213k processed examples, four types of tasks were generated (captioning, multiple-choice QA, open-ended QA, creative writing), resulting in approximately 800k-900k final examples depending on the inclusion of semantic descriptors (see Appendix B.3 for details). When comparing against other datasets, there was an overlap of only 12 examples with AudioSet, with one example also being in AudioCaps, and 0 in VGGSound.
@@
Both configurations use XML-style tags to distinguish between phases. In the two-phase approach, the model’s output is structured with <thinking> and <answer> tags. The three-phase approach introduces an additional <semantic_elements> section between thinking and answering, allowing the model to explicitly represent key semantic descriptors before generating the final caption. See also Appendix D with Figure 5 for an example of an output from the two-phase approach, and Figure 6 for an example of an output from the three-phase approach.
@@
A small subset on 100 randomly selected examples from the dataset was systematically evaluated with three annotators to assess both correctness and depth of responses on a Likert scale from 1 to 5. For correctness of the captions in terms of how well they describe the events happening in the audio, we achieved moderate to substantial agreement across multiple metrics: Krippendorff’s Alpha of 0.538, ICC(2,1) of 0.589, and weighted Kappa values ranging from 0.464 to 0.529 using linear weighting. The rating distributions show strong consistency, with between 70-77% of samples receiving the highest rating (5/5) across all three annotators. For depth of responses, while agreement was lower (Krippendorff’s Alpha of 0.207), the ratings remained consistently high with almost all samples (85%+) receiving the highest score.
@@ 3.1 Semantic Descriptors
The core idea behind incorporating semantic descriptors into model reasoning is to capture the detailed aspects of audio that a human listener would naturally perceive (see also [23, 62]). Before generating a final caption, the model first engages in a structured “thinking” phase, during which it attempts to identify and describe specific components of the audio signal. This process enables the model to move beyond surface-level recognition toward a more nuanced understanding of what it is “hearing”. The descriptors cover both objective and subjective dimensions of auditory perception. Objective descriptors include elements that produce the sound, such as sound-generating agents and physical sound sources, while subjective descriptors address how the sound is perceived, including auditory attributes and non-auditory sensations. This structured reasoning approach allows the model to decompose complex audio scenes into their core semantic descriptors, resulting in more informative and interpretable outputs. See Table 2 for definitions of the semantic descriptors used in our approach.
@@ 4 AudSemThinker Model
The proposed approach employs the Qwen2.5-Omni [67] model with 7B parameters. Qwen2.5-Omni is an end-to-end multimodal model that processes diverse inputs including text, images, audio, and video, and generates both text and speech outputs through its Thinker-Talker architecture. For AudSemThinker, the “Thinker” component, the large language model responsible for text generation, is specifically fine-tuned. This architecture processes multi-modal inputs and generates textual outputs. Fine-tuning occurs exclusively on the AudSem dataset (see Section 3); no other fine-tuning is performed. The core idea involves guiding the model to first engage in a “thinking” process, describe the semantic descriptors of the audio, and then produce a final “answer”. Two primary training paradigms are explored: SFT and GRPO, both using the AudSem dataset, demonstrating the dataset’s effectiveness in both scenarios.
@@ 4.1 Supervised Fine-Tuning (SFT)
Parameter-efficient fine-tuning is used with LoRA [30], targeting projection layers. Both models are trained using the AdamW optimizer [43] with a learning rate of 2e-04 for one epoch, employing a scheduler with linear decay and bf16 precision. Training uses a batch size of four on a single H100 GPU, taking approximately 12 hours for the full dataset (AudSemThinker) and six hours for the QA subset (AudSemThinker-QA). A detailed instructional prompt is prepended to the user’s question (see Appendix E.3), where the model is guided on the expected output format and reasoning depth (see Appendix 5.3 for a detailed ablation study).
@@ 4.2 Reinforcement Learning with Group Relative Policy Optimization (GRPO)
Beyond supervised fine-tuning, we investigate the applicability of reinforcement learning to our task and dataset using GRPO [55]. GRPO has shown to be an effective technique in finetuning models and has been used to train state-of-the-art models such as DeepSeek R1 [55] and Qwen2.5 [67]. GRPO allows us to explore whether the AudSem dataset is suitable for RL-based finetuning and to specifically test the impact of controlling the model’s “thinking budget” during generation. GRPO with Verifiable Rewards (RLVR) [9, 56] is employed, building upon a Qwen2.5-Omni base model.
@@
  • Length Constraint Reward: Specifically targeting the <think> phase, this reward function implements a “thinking budget”. It penalizes deviations from a target length for the reasoning content, using parameters to control the penalty strength and tolerance. We adapt the approach from [2] (with $\alpha=0.1,\delta=0.5$) by adding a reward component for being close to the target length (see Appendix C for details). This encourages the model to generate thinking content near a specified target length while penalizing excessive length.
These reward functions produce scalars for each generation, which are weighted and summed. For the group of generations corresponding to each prompt, the mean and standard deviation of these summed rewards are calculated. An advantage (summed reward - group mean) is computed for each generation, and these advantages are optionally normalized using the group’s standard deviation to produce the final scalar reward signal used for policy updates.
@@
The resulting model, AudSemThinker-QA GRPO, is trained with a target thinking budget of 25 words. Similar to SFT, LoRA [30] is used for parameter-efficient tuning. The input prompts for GRPO also include instructions referencing the target thinking length. For training, we use the default GRPO loss type with $\beta=0.01$, generate six completions per prompt ($k=6$), and employ a batch size of two per device with bf16 precision. The model is trained on four H100 GPUs for approximately 10 hours, utilizing DeepSpeed ZeRO-3 [53] and vLLM [36] for efficient training and inference.